OpenStudy (anonymous):

Evaluate the integral 1/sqrt[1-x^2], x= 1/2 to sqrt(3)/2

OpenStudy (anonymous):

\[\int\limits_{1/2}^{\sqrt{3}/2} 1/\sqrt{1-x^2} d\]

OpenStudy (anonymous):

arcsin(x). That's a inverse trig. integral.

OpenStudy (anonymous):

Then evaluate, of course :-)

OpenStudy (anonymous):

let u=arcsin(x), dv=dx, ,, find du..and v..

OpenStudy (anonymous):

ty bmp- i woulda go for trig sub... would be a mess lol-

OpenStudy (anonymous):

u would or wouldnt use inverse trig.??

OpenStudy (anonymous):

i would if bmp didnt mention that.== it lot easier to use arcsin(x)

OpenStudy (anonymous):

@sam30317 You should remember some main integrals: arctan, arcsin, arccos (albeit this is rare), to avoid doing extra work. :-) @iHelp Sure thing, :-). These kind of problem are generally to catch people offguard.

OpenStudy (anonymous):

This* kind of problems, damn my "grammer" is bad today.

OpenStudy (anonymous):

nvm that quest

OpenStudy (anonymous):

rofl, i would thought that was correct grammer--- i HATE ENGLISH =(

OpenStudy (anonymous):

I thought i did memorize them!!! Teachers never really give u examples or quiz u on this stuff. But on the test, they;re everywhere! would the answer be pi/6?

OpenStudy (anonymous):

no offense but 10 year or English, i still cant learn proper grammar.. it a pain

OpenStudy (anonymous):

Americans and Brits dont know their english that well either. Dont worry, its idiotic to us too

OpenStudy (anonymous):

Yup, that's correct @sam30317 . @iHelp None taken, I am not a native English speaker, haha.

OpenStudy (anonymous):

**Gasp** Omg, really??? Muah ha ha ha!

OpenStudy (anonymous):

btw sam, i have faith in u =) dont let us down

OpenStudy (anonymous):

lmao!! I'd rather you have faith in your deity and pray for me. I need allll the help i can get

OpenStudy (anonymous):

ill pray for us*

OpenStudy (anonymous):

Haha, lets do that. Ive been praying for weeks now.

OpenStudy (anonymous):

That's right. i have 3 in one day. I WILL RAPE THOSE PAPERS DOWN. amen.

OpenStudy (anonymous):

Omg, thats hilarious, lol!!!!! I need to get off here, haha